OpenStudy (anonymous):

Vector B (with arrow above) has x, y, and z components of 6.00, 8.00, and 5.00 units, respectively. Calculate the magnitude of B (with arrow above). ______ Calulcate the angle that makes with the x axis. ____° Calculate the angle that makes with the y axis. ____° Calculate the angle that makes with the z axis. ____°

OpenStudy (anonymous):

√(6^2+8^2+5^2) = √(36+64+25) = √125

OpenStudy (anonymous):

i got that part so far

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

I would draw a diagram and use trig to find the angles.

OpenStudy (anonymous):

like a triangle?

OpenStudy (anonymous):

Yeah, using the point you were given and then use trig (inverse tangent) to find the angles

OpenStudy (anonymous):

so x would arc tan of (5/8) y = arc tan of (6/5) z= arc tan of 6/8 ??

OpenStudy (anonymous):

For x use x,y For y use y,z For z use x,z

OpenStudy (anonymous):

so then 6/8,8/5,and 6/5?

OpenStudy (anonymous):

sorry =( i just want to get this right =(

OpenStudy (anonymous):

ill try it

OpenStudy (anonymous):

nope =(

OpenStudy (anonymous):

anyone?